Time for a lump sum to double

If you deposit money today in an account that pays 15% annual interest, how long will it take to double your money? Round your answer to two decimal places.

 years
Answer #1

Assume you deposit amount X today

Hence its future value = 2X

Using formula for future value of sum , we calculate the period required to double the money

Future value of ordinary sum = P * (1+r)^n

r = rate of interest = 15% per annum

n = no.of years

2X = X * (1+0.15)^n

2 = 1.15^n

n = 4.96

It will take 4.96 years to double your money.


answered by: FunkFuzzy